OpenStudy (anonymous):

How do i add or subtract Mixed numbers? im homeschooled, and don't have a teacher to ask ! ):

OpenStudy (anonymous):

what do you mean by mixed numbers?

OpenStudy (anonymous):

There like i guess, for an example 1 1/5 - 7 1/7

OpenStudy (anonymous):

as in 1 and 1/5?

OpenStudy (anonymous):

arrgg.

OpenStudy (anonymous):

when you have a mixed number you can think of it as addition so for example 2 4/5=2+4/5 so if you want to add 2 4/5 and 4 3/5 than if you rewrite both as addition it will make things easier 2+4/5+4+3/5=6+7/5=7+2/5= 7 2/5

OpenStudy (anonymous):

I never liked using mixed numbers, do not really see the point...

OpenStudy (anonymous):

so you'd rewrite it as 6/5 because 1 is the same as 5/5 and you'd rewrite 7 and 1/7 as 50/7 because 7 is the same as 49/7 then you'd make the denominator the same in this case I'd pick 35 so 6/5 is the same as 42/35 and 50/7 is the same as 250/35 then 42/35 - 250/35 = (42-250)/35=-208/35 which written as a mixed number is -5 and 33/35

OpenStudy (anonymous):

I prefer mixed numbers it makes it easier for me to visualise the amount

OpenStudy (anonymous):

true but to calculate with them is more work

OpenStudy (anonymous):

agreed